# On the Compatibility of Constructive Predicative Mathematics with Weyl’s Classical Predicativity

### v1

@@ 3.3 Homotopy Type Theory
with $A,B$ h-sets and $R(x,y)$ h-proposition for $x:A,y:B$, as shown in [53]. Indeed, in HoTT it is only possible to derive the axom of unique choice AC! formulated as follows
@@
In any case, it must be noticed that in HoTT real numbers a’ la Cauchy are defined as a higher inductive type $\mathbb{R}_{\mathsf{c}}$ ([53], see Sec. 11.3). Hence, they form a set. However, the construction is different from the usual type-theoretic definition of Cauchy reals or Bishop reals. In particular, the definition of $\mathbb{R}_{\mathsf{c}}$ is intended to formalize the idea that Cauchy reals can be regarded as the free complete metric space generated by the rationals.
@@
In this section, we want to show that already the weaker system of Heyting Arithmetic in all finite types (${\bf HA^{\omega}}$), when extended with the internal rule of unique choice and the law of excluded middle, becomes impredicative. Since the most common constructive and predicative foundations interpret this system, then the result can be easily extended to these more expressive theories.
@@
The language of ${\bf HA^{\omega}}$ includes countably infinite variables $x^{\sigma},y^{\sigma},z^{\sigma},\ldots$ for each type symbol $\sigma\in\mathrm{T}$. A binary predicate for equality $=_{\sigma}$ at each type $\sigma\in\mathrm{T}$ and an application operator $\mathsf{Ap}_{\sigma,\tau}$ for all $\sigma,\tau\in\mathrm{T}$. Furthermore, the language comprehends constants for zero and successor, recursor, pairing and projections and the combinators ${\bf k}$ and ${\bf s}$.
@@
We recall two fundamental results about ${\bf HA^{\omega}}$. It is immediate to realize that Heyting Arithmetic (HA) is a subsystem of ${\bf HA^{\omega}}$. Furthermore, if we extend ${\bf HA^{\omega}}$ with the Axiom of Choice for all finite types
$({\bf AC}_{\sigma,\tau})\ \forall x^{\sigma}\ \exists y^{\tau}\ \phi(x,y)\ \supset\ \exists f^{\sigma\rightarrow\tau}\ \forall x^{\sigma}\ \phi(x,fx)$
@@
Another significant fact is that ${\bf HA^{\omega}}$ is equi-consistent with its classical version by a double negation translation. This result can be found in [51]. However, this result does not apply to the case of ${\bf HA^{\omega}}+{\bf AC}_{\sigma,\tau}$. Indeed, it is possible to show that ${\bf HA^{\omega}}+{\bf AC}_{\sigma,\tau}+{\bf LEM}$ is enough strong to derive the second-order comprehension axiom and thus to interpret full second-order classical arithmetic ${\bf PA^{2}}$, which is an impredicative theory. The first to observe that ${\bf HA^{\omega}}$ extended with the Axiom of Choice and classical logic allows to derive the second-order comprehension principle was Spector in [49]. We remark that already the internal rule of unique choice (${\bf iRC!}_{\sigma,\tau}$) is sufficient to carry out the proof.
@@
The internal rule of unique choice for ${\bf HA^{\omega}}$ is formulated as follows:
@@
Therefore, we can apply ${\bf iRC!}_{\sigma,\tau}$ (for $\sigma,\tau=\mathbf{N}$) and we obtain that
@@ 5 The incompatibility of MLTT, HoTT and CZF with Weyl’s predicativity
Therefore, when MLTT is extended with the law of excluded middle, we can reproduce the argument showing the incompatibility of HAω with Weyl’s predicativity:
@@
MLTT with the addition of the law of excluded middle becomes impredicative.
@@
As in the case of MLTT, we can reproduce the same proof of Prop.4.3 but with a significant difference. The law of excluded middle cannot be formulated according to the propositions-as-types identifications for all types in HoTT, because this would yield a contradiction in presence of Univalence (see Ch.3 [53]). However, it is still possible to formulate ${\bf LEM}$ for all those types that are h-propositions. This formulation serves our purposes and is consistent with Univalence.
@@ 6 A possible way out: the Minimalist Foundation
The set-constructors of mTT and emTT  include those of first order Martin-Löf’s type theory, respectively as presented in [39] and [38]. Hence we have the following set constructors: the empty set $N_{0}$, the singleton set $N_{1}$, the set of lists $List(A)$ over a set $A$, hence the set of natural numbers $\mathsf{N}$, the indexed sum and the dependent product of the family of sets $B(x)\ set\ [x\in A]$ denoted respectively as $\Sigma_{x\in A}\ B(x)$ and $\Pi_{x\in A}\ B(x)$, the disjoint sum $A+B$ of the set $A$ with the set $B$. In the case of the extensional level emTT  we have that sets are closed under effective quotients $A/R$ over a set $A$, provided that $R$ is a small equivalence relation $R(x,y)\ prop_{s}\ [x\in A,y\in A]$. This set constructors is not included in the intensional level mTT. Furthermore, both the sets of mTT and those of emTT include also their small propositions $\phi\ prop_{s}$ regarded as sets of their proofs. Small propositions are those propositions closed under intuitionistic connectives and under quantifiers and equalities restricted to sets.
Collections of mTT and emTT include their sets and the indexed sum $\Sigma_{x\in A}B(x)$ of the family of collections $B(x)\ coll\ [x\in A]$ indexed over a collection $A$. Moreover, they also include propositions $\psi\ prop$ regarded as collections of their proofs. A significant distinction between mTT-collections and emTT-collections is the following: the former includes the proper collection of small propositions $\mathsf{prop_{s}}$ and the collection of small propositional functions $A\rightarrow\mathsf{prop_{s}}$ over a set $A$ (which are definitely not sets predicatively when $A$ is not empty), while the latter includes the power-collection of the singleton ${\cal P}(1)$, which is the quotient of the collection of small propositions under the relation of equiprovability, and the power-collection $A\rightarrow{\cal P}(1)$ of a set $A$, that can be written simply as ${\cal P}(A)$.
In mTT  we identify predicates over the set $A$ with propositional functions defined as primitive $\lambda$-terms. Hence, in emTT a subset of a set $A$ is identified with the equivalence class of propositional function up to equiprovability. So this use of $\lambda$-terms corresponding to predicates is a major difference with axiomatic set theory where subsets of a set $A$ are simply defined as sets in a certain relation with $A$ and hence are identified with functional relations into the boolean set $\{0,1\}$ when the logic is classical.
@@
It is relevant for our discussion to remark that elimination rules for propositions act only towards propositions and not towards collections. The same applies to small propositions. In this way, mTT and emTT do not generally validate choice principles. In particular, even unique choice cannot be derived in MF, thanks to the results in [30, 31, 25] and similarly to what happens in Coquand’s Calculus of Constructions CoC, as first proved in [50]. This is in contrast with what happens in MLTT and in HoTT, as we have seen in Section 3. The fact that choice principles are not validated in MF is important in connection with the notion of compatibility introduced in Section 5. In this way, MF turns out to be also compatible with theories where choice does not generally hold, such as CoC.
@@
From [25] we know that the rule of unique choice is not valid in mTT:
@@
The existence property of mTT, following from normalization results, would allow to deduce the equivalence of the rule of unique choice with its internal version and hence the non-validity of the internal rule of unique choice in mTT, too.
Nevertheless, it is worth pointing out that without relying on the existence property, we can show the not validity of the internal rule of unique choice following the same argument in [25] since this can be exported also to extensions of mTT, not validating the existence property, such as its classical version:
@@
It is important to remark that already the intensional level mTT can be regarded as an extension of ${\bf HA}^{\omega}$. Indeed, it is possible to show that there is a compatible translation of this theory into mTT following a similar translation of Section 5 for MLTT by interpreting ${\bf HA}^{\omega}$-propositions as mTT-propositions. What is most interesting is that MF  extends ${\bf HA}^{\omega}$ while preserving the fundamental property of equi-consistency with its classical counterpart. Indeed, recently it has been shown that both levels of MF are equi-consistent with their classical counterparts (collapsed to the extension of the extensional level emTT with the law of excluded middle), unlike theories such as MLTT, HoTT and CZF:
@@
Both levels of MF are equi-consistent with emTT with the addition of the law of excluded middle.
@@
For a proof, we refer to [27, 44]. ∎
Therefore, the extension of MF with the law of excluded middle is still predicative. From this, knowing that HAω can be embedding in mTT and emTT in a compatible way, we can deduce the non-derivability of unique choice on natural numbers in their classical version:
@@
The classical version of emTT and of mTT do not validate the number-theoretic internal rule of choice and hence the number-theoretic axiom of unique choice:
@@
First, observe that $\bf AC!_{\mathbb{N},\mathbb{N}}$ implies the number-theoretic internal rule of unique choice. Then, the claim follows from Prop. 6.5 and Cor. 4.4. ∎
Since the validity of the axiom of unique choice in a theory can be transferred to its extensions, contrary to the rule of unique choice, from Cor. 6.6 we can immediately deduce the non-validity of ${\bf AC!_{\mathbb{N},\mathbb{N}}}$ in mTT and emTT themselves:
@@
The intuitionistic version of ACA is interpretable in the extensional level emTT  of MF. Hence, ACA can be interpreted in its classical version.
@@
Peano axioms can be derived from the rules for natural numbers in emTT and the presence of power-collections, see [24]. The induction scheme of ACA follows from the elimination rule for $\mathsf{N}$ which acts towards all types including all propositions. Finally, the arithmetical comprehension axiom is validated as follows: given an arithmetical formula $\phi(x)$ in ACA, we know that it is interpreted as a small proposition in emTT. Then we can introduce $[\phi](x)\in[x\in\mathsf{N}]$ in $\mathcal{P}(1)$ and we get $X\ {\ :\equiv\ }\lambda x.\phi(x)\in\mathsf{N}\rightarrow\mathcal{P}(1)$. Therefore, it follows that there exists $X$ such that for all $n\in\mathsf{N}$, $X(x)\leftrightarrow\phi$. ∎
Hence, from prop. 6.8 it follows that the classical version of MF can formalize Weyl’s classical predicative reformulation of analysis in Das Kontinuum following Feferman’s analysis in [17, 18].
Furthermore, from [24, 20] we know a proof-theoretic upper bound for MF:
Proposition 6.9.
Corollary 6.10.
Proof.
It follows from Prop.6.5 and Prop.6.9.
@@
A direct proof of this was given [26] but it follows also from the predicativity of the classical version of MF. If real numbers form a set in MF then this is so in the classical version of MF and every interval of real numbers form a set too. Since classically, the power-collection of natural numbers can be put in bijection with an interval of real numbers, then it would follow that impredicative definitions can be employed in the classical version of MF contrary to its predicativity established thanks in prop.6.10. ∎
@@ 7 Conclusion
Given that the Minimalist Foundation interprets HAω, the price to pay is to renounce the rule of unique choice as well as all the number-theoretic choice principles characteristic of Bishop’s mathematics and furthermore that real numbers à la Dedekind or Cauchy form a set. All these facts together call for a point-free development of analysis by adopting the topological methods advocated by Martin-Löf and Sambin with the introduction of Formal Topology [45].
@@
In the future, besides establishing the exact proof-theoretic strength of MF, which is currently an open problem, we will explore whether the extension of MF  with inductive and coinductive definitions in [28, 29, 32] is still predicative, or even is still equiconsistent with its classical counterpart, as established for MF in [27, 44]. A positive answer to either of these questions would provide a foundation capable of formalizing general results of point-free mathematics as those presented in [12, 36, 15, 47], and still compatible with both Bishop’s constructivism and Weyl’s classical predicativity, as in the case of MF.

### v2

@@ 3.3 Homotopy Type Theory
with $A,B$ h-sets and $R(x,y)$ h-proposition for $x:A,y:B$, as shown in [53]. Indeed, in HoTT it is only possible to derive the axiom of unique choice AC! formulated as follows
@@
In any case, it must be noticed that real numbers à la Cauchy in HoTT are defined in a different way from the usual type-theoretic definition of Cauchy reals or Bishop reals. Indeed, they are defined as a higher inductive type $\mathbb{R}_{\mathsf{c}}$ ([53], see Sec. 11.3) capturing the idea that Cauchy reals can be regarded as the free complete metric space generated by the rationals and they form an h-set.
@@
In this section, we want to show that already the weaker system of Heyting Arithmetic in all finite types (${\bf HA^{\omega}}$), when extended with the number-theoretic internal rule of unique choice and the law of excluded middle, becomes impredicative. Since the most common constructive and predicative foundations interpret this system, then the result can be easily extended to these more expressive theories.
@@
The language of ${\bf HA^{\omega}}$ includes countably infinite variables $x^{\sigma},y^{\sigma},z^{\sigma},\ldots$ for each type symbol $\sigma\in\mathrm{T}$. A binary predicate for equality $=_{\sigma}$ at each type $\sigma\in\mathrm{T}$ and an application operator $\mathsf{Ap}_{\sigma,\tau}$ for all $\sigma,\tau\in\mathrm{T}$. Furthermore, the language comprehends constants for zero and successor, recursor, pairing and projections and the combinators ${\bf k}$ and ${\bf s}$. Note that the logical implication is denoted with the symbol $\supset$ to avoid confusion with the symbol $\rightarrow$ used to denote function types.
@@
We recall two fundamental results about ${\bf HA^{\omega}}$. It is immediate to realize that Heyting Arithmetic (HA) is a subsystem of ${\bf HA^{\omega}}$. Furthermore, if we extend ${\bf HA^{\omega}}$ with the Axiom of Choice for all finite types $\sigma,\tau$:
$({\bf AC})\ \ \forall x^{\sigma}\ \exists y^{\tau}\ \phi(x,y)\ \supset\ \exists f^{\sigma\rightarrow\tau}\ \forall x^{\sigma}\ \phi(x,fx)$
@@
Another significant fact is that ${\bf HA^{\omega}}$ is equiconsistent with its classical version by a double negation translation. This result can be found in [51]. However, this result does not extend to the case of ${\bf HA^{\omega}}+{\bf AC}$. Indeed, it is possible to show that ${\bf HA^{\omega}}+{\bf AC}+{\bf LEM}$, where LEM is the law of excluded middle $\phi\vee\neg\phi$ for all formulas, is strong enough to derive the second-order comprehension axiom and thus to interpret full second-order classical arithmetic ${\bf PA^{2}}$, which is an impredicative theory. The first to observe that ${\bf HA^{\omega}}$ extended with the Axiom of Choice and classical logic allows one to derive the second-order comprehension principle was Spector in [49]. We remark that already the internal number-theoretic rule of unique choice is sufficient to carry out the proof.
@@
The internal rule of unique choice, called ${\bf iRC!}$, is formulated for ${\bf HA^{\omega}}$ as follows:
@@
The special case when ${\bf iRC!}$ acts only on natural numbers, namely both $\sigma$ and $\tau$ coincide with the type $\mathbf{N}$ of natural numbers, is called internal number-theoretic rule of unique choice and it is denoted with the symbol ${\bf iRC!_{\mathbf{N},\mathbf{N}}}$.
@@
Therefore, we can apply ${\bf iRC!}_{\mathbf{N},\mathbf{N}}$ and we obtain that
@@ 5 The incompatibility of MLTT, HoTT and CZF with Weyl’s predicativity
Therefore, suppose to extend MLTT with the law of excluded middle LEM formulated as the existence of a proof-term of the type
$A+(A\rightarrow N_{0})$
where $A$ is any type, $N_{o}$ is the empty set interpreting the falsum constant, $+$ is the symbol for the disjoint sum type and $\rightarrow$ is the symbol for the function type. Then in MLTT+LEM we can reproduce the argument showing the incompatibility of HAω with Weyl’s predicativity:
@@
MLTT + LEM becomes impredicative.
@@
As in the case of MLTT, we can reproduce the same proof of Prop.4.3 but with a significant difference. The law of excluded middle cannot be formulated according to the propositions-as-types identifications for all types in HoTT, because this would yield a contradiction in presence of Univalence (see Ch.3 [53]). However, it is still possible to formulate ${\bf LEM}$ for all those types that are h-propositions as
$\phi\vee\neg\phi$
for all h-propositions $\phi$.
This formulation serves our purposes and is consistent with Univalence.
@@ 6 A possible way out: the Minimalist Foundation
The set-constructors of mTT and emTT  include those of first order Martin-Löf’s type theory, respectively as presented in [40] and [39]. Hence we have the following set constructors: the empty set $N_{0}$, the singleton set $N_{1}$, the set of lists $List(A)$ over a set $A$, hence the set of natural numbers $\mathsf{N}$, the indexed sum and the dependent product of the family of sets $B(x)\ set\ [x\in A]$ denoted respectively as $\Sigma_{x\in A}\ B(x)$ and $\Pi_{x\in A}\ B(x)$, the disjoint sum $A+B$ of the set $A$ with the set $B$. In the case of the extensional level emTT  we have that sets are closed under effective quotients $A/R$ over a set $A$, provided that $R$ is a small equivalence relation $R(x,y)\ prop_{s}\ [x\in A,y\in A]$. Such quotient constructors are not included in the intensional level mTT. Furthermore, both mTT and emTT include small propositions $\phi\ prop_{s}$ regarded as sets of their proofs and defined are those propositions closed under intuitionistic connectives, quantifiers, and equalities restricted to sets.
Collections of mTT and emTT include their sets, the indexed sum $\Sigma_{x\in A}B(x)$ of the family of collections $B(x)\ coll\ [x\in A]$ indexed over a collection $A$, and propositions $\psi\ prop$ regarded as collections of their proofs. A significant distinction between mTT-collections and emTT-collections is the following: the former includes the proper collection of small propositions $\mathsf{prop_{s}}$ and the collection of small propositional functions $A\rightarrow\mathsf{prop_{s}}$ over a set $A$ (which are never sets predicatively when $A$ is not empty), while the latter includes the power-collection of the singleton ${\cal P}(1)$, which is the quotient of the collection of small propositions under the relation of equiprovability, and the power-collection $A\rightarrow{\cal P}(1)$ of a set $A$, that can be written simply as ${\cal P}(A)$.
In mTT  we identify predicates over the set $A$ with propositional functions defined as primitive $\lambda$-terms. Hence, in emTT a subset of a set $A$ is identified with the equivalence class of propositional function up to equiprovability. So this use of $\lambda$-terms corresponding to predicates is a major difference with an axiomatic set theory where subsets of a set $A$ are simply defined as sets in a certain relation with $A$ and hence are identified with functional relations into the boolean set $\{0,1\}$ when the logic is classical.
@@
It is relevant for our discussion to remark that elimination rules for propositions act only towards propositions and not towards collections. The same applies to small propositions. In this way, mTT and emTT do not generally validate choice principles. In particular, even unique choice cannot be derived in both levels of MF, thanks to results in [30, 31, 25] and the non-derivability of unique choice in Coquand’s Calculus of Constructions CoC, as first proved in [50]. This is in contrast with what happens in MLTT and in HoTT, as we have seen in Section 3. The fact that choice principles are not validated in MF is important in connection with the notion of compatibility introduced in Section 5. In this way, MF turns out to be also compatible with theories where choice does not generally hold, such as CoC.
@@
Concerning this rule observe that:
@@
The existence property of mTT would allow to deduce the equivalence of the rule of unique choice with its internal version and hence the non-validity of the internal rule of unique choice in mTT, too.
Nevertheless, it is worth pointing out that without relying on the existence property, we can show the non-validity of the internal rule of unique choice following the same argument in [25] since this can also be exported to extensions of mTT, not validating the existence property, such as its classical version:
@@
It is important to remark that already the intensional level mTT can be regarded as an extension of ${\bf HA}^{\omega}$. Indeed, it is possible to show that there is a compatible translation of this theory into mTT following a similar translation of Section 5 for MLTT by interpreting ${\bf HA}^{\omega}$-propositions as mTT-propositions. What is most interesting is that MF  extends ${\bf HA}^{\omega}$ while preserving the fundamental property of equi-consistency with its classical counterpart. Indeed, recently it has been shown that both levels of MF are equiconsistent with their classical counterparts (collapsed to the extension of the extensional level emTT with the law of excluded middle), unlike theories such as MLTT, HoTT and CZF:
@@
Both levels of MF are equiconsistent with emTT extended the addition of the law of excluded middle $\phi\vee\neg\phi$ for all propositions $\phi$.
@@
For a proof, we refer to [33]. ∎
Furthermore, from [24, 20] we know a proof-theoretic upper bound for MF:
Proposition 6.6.
Therefore, the extension of MF with the law of excluded middle is still predicative:
@@
Proof.
It follows from Prop.6.5 and Prop.6.6.
From this, knowing that HAω can be embedded in mTT and emTT in a compatible way, we can deduce the non-derivability of unique choice on natural numbers in their classical version:
Corollary 6.8.
The classical versions of emTT and mTT do not validate the number-theoretic internal rule of choice and hence the number-theoretic axiom of unique choice:
@@
First, observe that $\bf AC!_{\mathbb{N},\mathbb{N}}$ implies the number-theoretic internal rule of unique choice. The claim then follows from Proposition 6.7 and Cor. 4.4. ∎
Since the validity of the axiom of unique choice in a theory can be transferred to its extensions, contrary to the rule of unique choice, from Cor. 6.8 we can immediately deduce the non-validity of ${\bf AC!_{\mathbb{N},\mathbb{N}}}$ in mTT and emTT themselves:
@@
The intuitionistic version of ACA is interpretable in the extensional level emTT  of MF. Hence, (the classical) ACA can be interpreted in the classical version of emTT.
@@
Peano axioms can be derived from the rules for natural numbers in emTT and the presence of power-collections, see [24]. The induction scheme of ACA follows from the elimination rule for $\mathsf{N}$ which acts towards all types including all propositions. Finally, the arithmetical comprehension axiom is validated as follows: given an arithmetical formula $\phi(x)$ in ACA, we know that it is interpreted as a small proposition in emTT. Then we can introduce $[\phi](x)\in[x\in\mathbb{N}]$ in $\mathcal{P}(1)$ and we get $X\ {\ :\equiv\ }\lambda x.\phi(x)\in\mathbb{N}\rightarrow\mathcal{P}(1)$. Therefore, it follows that there exists $X$ such that for all $n\in\mathbb{N}$, $X(x)\leftrightarrow\phi$. ∎
Hence, from prop. 6.10 it follows that the classical version of MF can formalize Weyl’s classical predicative reformulation of analysis in Das Kontinuum following Feferman’s analysis in [17, 18].
@@
A direct proof of this was given [26] but it follows also from the predicativity of the classical version of MF (see also [33]). If real numbers form a set in MF then this is so in the classical version of MF and every interval of real numbers form a set too. Since classically, the power-collection of natural numbers can be put in bijection with an interval of real numbers, then it would follow that impredicative definitions can be employed in the classical version of MF contrary to its predicativity established thanks in prop.6.7. ∎
@@ 7 Conclusion
Given that the Minimalist Foundation interprets HAω, the price to pay is to renounce the rule of unique choice as well as all the number-theoretic choice principles characteristic of Bishop’s mathematics and that real numbers à la Dedekind or Cauchy form a set. All these facts together call for a point-free development of analysis by adopting the topological methods advocated by Martin-Löf and Sambin with the introduction of Formal Topology [45].
@@
In the future, besides establishing the exact proof-theoretic strength of MF, which is currently an open problem, we will explore whether the extensions of MF  with inductive and coinductive definitions in [28, 29, 32] are still equiconsistent with their classical counterpart, as established for MF in [33], or at least they are still predicative after the addition of the law of excluded middle. A positive answer to either of these questions would provide a more expressive constructive foundation capable of formalizing general results of point-free mathematics as those presented in [12, 37, 15, 47], by keeping the compatibility both with Bishop’s constructivism and with Weyl’s classical predicativity, as in the case of MF.